OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = (4(4n+1)(8n+7))/6 I understand how to prove it by mathematical induction for the most part, I just don't understand how the sequence can possibly go from multiplying to adding and back again.

OpenStudy (amistre64):

(4 ⋅ 6) + (5 ⋅ 7) + (6 ⋅ 8) + ... + 4n( 4n + 2) 4(4+2) + 5(5+2)+ 6(6+2)+ ... + 4n( 4n + 2)

OpenStudy (anonymous):

Oh gosh. That makes sense now. There was nothing pointing to that through all the material I reviewed.

OpenStudy (amistre64):

4(4+2) = (4(4+1)(8+7))/6 24 = 50 is not trtue

OpenStudy (anonymous):

Yes. This makes much more sense now. Thank you!